### v1

@@ Abstract
Recent advances in multi-modal detection have significantly improved detection accuracy in challenging environments (e.g., low light, overexposure). By integrating RGB with modalities such as thermal and depth, multi-modal fusion increases data redundancy and system robustness. However, significant challenges remain in effectively extracting task-relevant information both within and across modalities, as well as in achieving precise cross-modal alignment. While CNNs excel at feature extraction, they are limited by constrained receptive fields, strong inductive biases, and difficulty in capturing long-range dependencies. Transformer-based models offer global context but suffer from quadratic computational complexity and are confined to pairwise correlation modeling. Mamba and other State Space Models (SSMs), on the other hand, are hindered by their sequential scanning mechanism, which flattens 2D spatial structures into 1D sequences, disrupting topological relationships and limiting the modeling of complex higher-order dependencies. To address these issues, we propose a multi-modal perception network based on hypergraph theory called M2I2HA. Our architecture includes an Intra-Hypergraph Enhancement module to capture global many-to-many high-order relationships within each modality, and an Inter-Hypergraph Fusion module to align, enhance, and fuse cross-modal features by bridging configuration and spatial gaps between data sources. We further introduce a M2-FullPAD module to enable adaptive multi-level fusion of multi-modal enhanced features within the network, meanwhile enhancing data distribution and flow across the architecture. Extensive object detection experiments on multiple public datasets against baselines demonstrate that M2I2HA achieves state-of-the-art performance in multi-modal object detection tasks. Our code and pre-trained models will be available at: https://github.com/WSYANGSX/machine_learning.
@@ 1 Introduction
To overcome these bottlenecks, we introduce a novel multi-modal detection network called M2I2HA that enables structured multi-modal reasoning via a hypergraph attention mechanism. The core innovation lies in the synergy of hypergraph attention and multi-modal fusion. The hypergraph attention dynamically constructs high-order hyperedges among image patches, enabling it to capture arbitrary-order element collaborations and achieve holistic modeling of global high-order dependencies. Unlike sequential modeling and traditional graph structures, a single edge in a hypergraph can connect multiple vertices, directly modeling high-order associations among multiple vertices. Existing studies [42, 9, 27, 8] have shown that using hypergraph to model multi-pixel high-order associations in visual tasks is of great value. M2I2HA achieves cross-modal high-order information modeling and fusion through two key modules: the Intra-Hypergraph Enhancement module efficiently extracts high-order features within a modality, enhancing the network’s ability to model global many-to-many high-order relationships, and achieving feature enhancement; the Intra-Hypergraph Fusion module aligns, enhances, and fuses cross-modal features. By utilizing a cross-hypergraph enhancement mechanism, the module models high-order relationships between correlated nodes in two modal feature spaces through hyperedges, and subsequently enhances multi-modal features using a gated attention mechanism. This enables adaptive attention allocation and ultimately improves perceptual performance. To further enhance adaptive data fusion, and optimize data flow within the network, we introduce an improved M2-FullPAD module. This module facilitates the adaptive fusion of multimodal features across multiple levels, significantly enhancing data distribution and fusion efficiency, alleviating data bottlenecks in deep networks, and improving the detection accuracy of small targets. Experiments on multiple public datasets show that M2I2HA achieves significant improvements in object detection accuracy and environmental perception ability compared to existing methods without increasing model parameters or sacrificing computational efficiency and inference speed, verifying the effectiveness and superiority of the proposed method. In summary, our contributions are as follows:
@@
3. We propose an enhanced M2-FullPAD module, enabling the network to adaptively conduct multi-level fusion of features across different stages. This module not only facilitates the adaptive fusion of multimodal features at multiple levels but also significantly improves data distribution and integration efficiency. Consequently, it alleviates data bottlenecks in deep networks and enhances the detection accuracy of small targets.
@@ 3 Methodology
In this section, we provide a comprehensive introduction to the proposed M2I2HA multi-modal object detection network. This section begins with a description of the preliminaries of Hypergraph. It is followed by an overview of the M2I2HA architecture. Then, we conduct a detailed discussion on its key components: the Intra-Hypergraph Enhancement Module, the Inter-Hypergraph Fusion Module, and the M2-FullPAD block.
Figure 3: Overall architecture of M2I2HA. M2I2HA adopts a YOLO-based architecture comprising a backbone, neck, and head. The backbone consists of two parallel feature extraction networks, two Intra-Hypergraph Enhancement modules, and an Inter-Hypergraph Fusion module, which respectively perform multi-modal feature extraction, intra-modal enhancement, and cross-modal fusion. The neck employs an FPN/PAN structure for multi-scale feature aggregation, while the head handles object detection. Additionally, the M2-FullPAD block optimizes multi-level feature information flow and enables dynamic fusion within the network.
@@ 3.2 Overall Architecture
The extracted features $\{\mathcal{F}^{rgb}_{3},\mathcal{F}^{rgb}_{4},\mathcal{F}^{rgb}_{5},\mathcal{F}^{ir}_{3},\mathcal{F}^{ir}_{4},\mathcal{F}^{ir}_{5}\}$ are fed into their respective Intra-Hypergraph Enhancement modules. This enables adaptive modeling of high-order correlations across scales and spatial positions within each modality. Enhanced features $\{\mathcal{H}^{rgb}_{3},\mathcal{H}^{rgb}_{4},\mathcal{H}^{rgb}_{5},\mathcal{H}^{ir}_{3},\mathcal{H}^{ir}_{4},\mathcal{H}^{ir}_{5}\}$ for each modality are obtained as a result. Then, enhanced features $\{\mathcal{H}^{rgb}_{5},\mathcal{H}^{ir}_{5}\}$ are then fed into the Inter-Hypergraph Fusion module, which performs adaptive modeling of correlations across modalities and conducts fusion, yielding the cross-modal correlation representation $\{\mathcal{C}_{3},\mathcal{C}_{4},\mathcal{C}_{5}\}$. Utilizing high-level features helps mitigate the impact of spatial misalignment in multi-modal data, thereby enhancing the effectiveness of cross-modal information modeling. At the same time, performing inter-modal interactions exclusively with high-level features significantly reduces computational overhead and improves the model’s real-time performance [31]. Subsequently, the M2-FullPAD module leverages three extracted features per modality along with intra- and inter-modal enhanced features to allocate the correlation-enhanced features to the connections between the backbone, neck, and head. This method retains low-level features for small object detection while enabling a more effective and adaptive flow of information. In the neck part of the M2I2HA network, we adopt the same structure as YOLOv13, which employs large-kernel depthwise-separable convolution (DSConv) as the core component to construct lightweight feature extraction blocks. The Neck part adopts the classic Path Aggregation Network (PAN) and Feature Pyramid Network (FPN) structure to achieve efficient fusion and enhancement of multi-scale features. Through the bidirectional paths from top to bottom and bottom to top, this structure can fully integrate shallow localization information and deep semantic information, thereby significantly improving the model’s detection ability for targets of different scales. The output feature maps from the neck are then passed to the detection head to perform multi-scale object detection.
@@ 3.4 Intre-Hypergraph Fusion Module
3.5 M2-FullPAD Block
We proposed the M2-FullPAD block for the distribution and dynamic fusion of features and augmented data at all levels within the network. The Data Processing Inequality (DPI) clearly states that during the transmission of signals, the data contained within cannot be increased through any physical processing operation [52]. That is, given a signal processing flow $S_{1}\rightarrow S_{2}\rightarrow...\rightarrow S_{n}$, the amount of data in the signal flow either remains unchanged or decreases. In detail, the mutual information (MI) between $S_{1}$ and $S_{n-1}$ is no less than the mutual information between $S_{1}$ and $S_{n}$, i.e.,
@@ 3.5 M2-FullPAD Block
M2-FullPAD collects the multi-scale feature maps $\{\mathcal{F}^{rgb}_{i},\mathcal{F}^{ir}_{i}\}$ extracted from the backbone network, the intra modality enhanced features $\{\mathcal{H}^{rgb}_{i},\mathcal{H}^{ir}_{i}\}$ output by the Intra-Hypergraph Enhancement module, and the inter modality enhanced features $\{\mathcal{C}_{3},\mathcal{C}_{4},\mathcal{C}_{5}\}$ output by the Inter Hypergraph Fusion module. It then redistributes these enhanced features to various positions throughout the neck pipeline via different pathways, as shown in Figure 3. Inside M2-FullPAD, a adaptive fusion mechanism is employed to enable dynamic and adaptive allocation and fusion of the network features. For an arbitrary feature map
@@
Where $\text{ModalFuseSE}(\cdot)$ denotes the feature-level channel attention fusion module, whose structure is similar to FuseSEBlock in Figure 4. $\alpha$, $\beta$ and $\gamma$ are learnable scalar parameters that dynamically adapt the contribution of correlation-enhanced features and original features during training. Based on M2-FullPAD module, correlation-enhanced features are effectively integrated into different stages of the pipeline, enabling the model to fully leverage the correlation information between RGB and thermal modalities, thereby enhancing its perception of complex scenes.
@@ 4.5 Evaluation on the VEDAI Dataset
COMO achieved the second-highest overall mAP score of 44.2% on the VEDAI dataset. In parallel, our YOLOv8s-based model demonstrated strong competitiveness, securing the second-best mAP@0.5 with 74.5% and the third-best overall mAP with 43.6%, closely approaching COMO’s performance. This indicates that our method is also highly competitive in the field of small object remote sensing. This remarkable performance can be attributed to our proposed Hypergraph-based multi-modal feature fusion architecture and the M2-FullPAD module. This design not only effectively captures and integrates enhanced modal features, but also enables the M2-FullPAD module to adaptively preserve low-level features that are essential for detecting small objects. As a result, our model consistently demonstrates strong representational capacity even when handling datasets with limited scale or tasks involving very small targets.
@@ 4.6 Ablation Studies
In this section, we conduct ablation studies on the core components of the proposed M2I2HA framework to systematically evaluate the individual contributions. We adopt YOLOv8s as the baseline model and investigate the impact of three key modules: the Intra-Hypergraph Enhancement Module (IHE), the Inter-Hypergraph Fusion Module (IHF), and the M2-FullPAD Block (M2FP). For single-modal detection, YOLOv8s is used directly, whereas for multi-modal detection, we follow COMO [31] by integrating a simple convolutional fusion module after the feature extraction backbones, forming our multi-modal baseline. All experiments are conducted on the DroneVehicle dataset.

### v2

@@ Abstract
Recent advances in multi-modal detection have significantly improved detection accuracy in challenging environments (e.g., low light, overexposure). By integrating RGB with modalities such as thermal and depth, multi-modal fusion increases data redundancy and system robustness. However, significant challenges remain in effectively extracting task-relevant information both within and across modalities, as well as in achieving precise cross-modal alignment. While CNNs excel at feature extraction, they are limited by constrained receptive fields, strong inductive biases, and difficulty in capturing long-range dependencies. Transformer-based models offer global context but suffer from quadratic computational complexity and are confined to pairwise correlation modeling. Mamba and other State Space Models (SSMs), on the other hand, are hindered by their sequential scanning mechanism, which flattens 2D spatial structures into 1D sequences, disrupting topological relationships and limiting the modeling of complex higher-order dependencies. To address these issues, we propose a multi-modal perception network based on hypergraph theory called M2I2HA. Our architecture includes an Intra-Hypergraph Enhancement module to capture global many-to-many high-order relationships within each modality, and an Inter-Hypergraph Fusion module to align, enhance, and fuse cross-modal features by bridging configuration and spatial gaps between data sources. We further introduce a M3FDFP module to enable adaptive multi-level fusion of multi-modal enhanced features within the network, meanwhile enhancing data distribution and flow across the architecture. Extensive object detection experiments on multiple public datasets against baselines demonstrate that M2I2HA achieves state-of-the-art performance in multi-modal object detection tasks. Our code and pre-trained models will be available at: https://github.com/WSYANGSX/machine_learning.
@@ 1 Introduction
To overcome these bottlenecks, we introduce a novel multi-modal detection network called M2I2HA that enables structured multi-modal reasoning via a hypergraph attention mechanism. The core innovation lies in the synergy of hypergraph attention and multi-modal fusion. The hypergraph attention dynamically constructs high-order hyperedges among image patches, enabling it to capture arbitrary-order element collaborations and achieve holistic modeling of global high-order dependencies. Unlike sequential modeling and traditional graph structures, a single edge in a hypergraph can connect multiple vertices, directly modeling high-order associations among multiple vertices. Existing studies [42, 9, 27, 8] have shown that using hypergraph to model multi-pixel high-order associations in visual tasks is of great value. M2I2HA achieves cross-modal high-order information modeling and fusion through two key modules: the Intra-Hypergraph Enhancement module efficiently extracts high-order features within a modality, enhancing the network’s ability to model global many-to-many high-order relationships, and achieving feature enhancement; the Intra-Hypergraph Fusion module aligns, enhances, and fuses cross-modal features. By utilizing a cross-hypergraph enhancement mechanism, the module models high-order relationships between correlated nodes in two modal feature spaces through hyperedges, and subsequently enhances multi-modal features using a gated attention mechanism. This enables adaptive attention allocation and ultimately improves perceptual performance. To further enhance adaptive data fusion, and optimize data flow within the network, we introduce an improved M3FDFP module. This module facilitates the adaptive fusion of multimodal features across multiple levels, significantly enhancing data distribution and fusion efficiency, alleviating data bottlenecks in deep networks, and improving the detection accuracy of small targets. Experiments on multiple public datasets show that M2I2HA achieves significant improvements in object detection accuracy and environmental perception ability compared to existing methods without increasing model parameters or sacrificing computational efficiency and inference speed, verifying the effectiveness and superiority of the proposed method. In summary, our contributions are as follows:
@@
3. We propose an enhanced M3FDFP module, enabling the network to adaptively conduct multi-level fusion of features across different stages. This module not only facilitates the adaptive fusion of multimodal features at multiple levels but also significantly improves data distribution and integration efficiency. Consequently, it alleviates data bottlenecks in deep networks and enhances the detection accuracy of small targets.
@@ 3 Methodology
In this section, we provide a comprehensive introduction to the proposed M2I2HA multi-modal object detection network. This section begins with a description of the preliminaries of Hypergraph. It is followed by an overview of the M2I2HA architecture. Then, we conduct a detailed discussion on its key components: the Intra-Hypergraph Enhancement Module, the Inter-Hypergraph Fusion Module, and the M3FDFP block.
Figure 3: Overall architecture of M2I2HA. M2I2HA adopts a YOLO-based architecture comprising a backbone, neck, and head. The backbone consists of two parallel feature extraction networks, two Intra-Hypergraph Enhancement modules, and an Inter-Hypergraph Fusion module, which respectively perform multi-modal feature extraction, intra-modal enhancement, and cross-modal fusion. The neck employs an FPN/PAN structure for multi-scale feature aggregation, while the head handles object detection. Additionally, the M3FDFP block optimizes multi-level feature information flow and enables dynamic fusion within the network.
@@ 3.2 Overall Architecture
The extracted features $\{\mathcal{F}^{rgb}_{3},\mathcal{F}^{rgb}_{4},\mathcal{F}^{rgb}_{5},\mathcal{F}^{ir}_{3},\mathcal{F}^{ir}_{4},\mathcal{F}^{ir}_{5}\}$ are fed into their respective Intra-Hypergraph Enhancement modules. This enables adaptive modeling of high-order correlations across scales and spatial positions within each modality. Enhanced features $\{\mathcal{H}^{rgb}_{3},\mathcal{H}^{rgb}_{4},\mathcal{H}^{rgb}_{5},\mathcal{H}^{ir}_{3},\mathcal{H}^{ir}_{4},\mathcal{H}^{ir}_{5}\}$ for each modality are obtained as a result. Then, enhanced features $\{\mathcal{H}^{rgb}_{5},\mathcal{H}^{ir}_{5}\}$ are then fed into the Inter-Hypergraph Fusion module, which performs adaptive modeling of correlations across modalities and conducts fusion, yielding the cross-modal correlation representation $\{\mathcal{C}_{3},\mathcal{C}_{4},\mathcal{C}_{5}\}$. Utilizing high-level features helps mitigate the impact of spatial misalignment in multi-modal data, thereby enhancing the effectiveness of cross-modal information modeling. At the same time, performing inter-modal interactions exclusively with high-level features significantly reduces computational overhead and improves the model’s real-time performance [31]. Subsequently, the M3FDFP module leverages three extracted features per modality along with intra- and inter-modal enhanced features to allocate the correlation-enhanced features to the connections between the backbone, neck, and head. This method retains low-level features for small object detection while enabling a more effective and adaptive flow of information. In the neck part of the M2I2HA network, we adopt the same structure as YOLOv13, which employs large-kernel depthwise-separable convolution (DSConv) as the core component to construct lightweight feature extraction blocks. The Neck part adopts the classic Path Aggregation Network (PAN) and Feature Pyramid Network (FPN) structure to achieve efficient fusion and enhancement of multi-scale features. Through the bidirectional paths from top to bottom and bottom to top, this structure can fully integrate shallow localization information and deep semantic information, thereby significantly improving the model’s detection ability for targets of different scales. The output feature maps from the neck are then passed to the detection head to perform multi-scale object detection.
@@ 3.4 Intre-Hypergraph Fusion Module
3.5 M3FDFP Block
We proposed the Multimodal and Multi-level Feature Dynamic Fusion Pipeline (M3FDFP) block for the distribution and dynamic fusion of features and augmented data at all levels within the network. The Data Processing Inequality (DPI) clearly states that during the transmission of signals, the data contained within cannot be increased through any physical processing operation [52]. That is, given a signal processing flow $S_{1}\rightarrow S_{2}\rightarrow...\rightarrow S_{n}$, the amount of data in the signal flow either remains unchanged or decreases. In detail, the mutual information (MI) between $S_{1}$ and $S_{n-1}$ is no less than the mutual information between $S_{1}$ and $S_{n}$, i.e.,
@@ 3.5 M3FDFP Block
M3FDFP collects the multi-scale feature maps $\{\mathcal{F}^{rgb}_{i},\mathcal{F}^{ir}_{i}\}$ extracted from the backbone network, the intra modality enhanced features $\{\mathcal{H}^{rgb}_{i},\mathcal{H}^{ir}_{i}\}$ output by the Intra-Hypergraph Enhancement module, and the inter modality enhanced features $\{\mathcal{C}_{3},\mathcal{C}_{4},\mathcal{C}_{5}\}$ output by the Inter Hypergraph Fusion module. It then redistributes these enhanced features to various positions throughout the neck pipeline via different pathways, as shown in Figure 3. Inside M3FDFP, a adaptive fusion mechanism is employed to enable dynamic and adaptive allocation and fusion of the network features. For an arbitrary feature map
@@
Where $\text{ModalFuseSE}(\cdot)$ denotes the feature-level channel attention fusion module, whose structure is similar to FuseSEBlock in Figure 4. $\alpha$, $\beta$ and $\gamma$ are learnable scalar parameters that dynamically adapt the contribution of correlation-enhanced features and original features during training. Based on M3FDFP module, correlation-enhanced features are effectively integrated into different stages of the pipeline, enabling the model to fully leverage the correlation information between RGB and thermal modalities, thereby enhancing its perception of complex scenes.
@@ 4.5 Evaluation on the VEDAI Dataset
COMO achieved the second-highest overall mAP score of 44.2% on the VEDAI dataset. In parallel, our YOLOv8s-based model demonstrated strong competitiveness, securing the second-best mAP@0.5 with 74.5% and the third-best overall mAP with 43.6%, closely approaching COMO’s performance. This indicates that our method is also highly competitive in the field of small object remote sensing. This remarkable performance can be attributed to our proposed Hypergraph-based multi-modal feature fusion architecture and the M3FDFP module. This design not only effectively captures and integrates enhanced modal features, but also enables the M3FDFP module to adaptively preserve low-level features that are essential for detecting small objects. As a result, our model consistently demonstrates strong representational capacity even when handling datasets with limited scale or tasks involving very small targets.
@@ 4.6 Ablation Studies
In this section, we conduct ablation studies on the core components of the proposed M2I2HA framework to systematically evaluate the individual contributions. We adopt YOLOv8s as the baseline model and investigate the impact of three key modules: the Intra-Hypergraph Enhancement Module (IHE), the Inter-Hypergraph Fusion Module (IHF), and the M3FDFP Block (M2FP). For single-modal detection, YOLOv8s is used directly, whereas for multi-modal detection, we follow COMO [31] by integrating a simple convolutional fusion module after the feature extraction backbones, forming our multi-modal baseline. All experiments are conducted on the DroneVehicle dataset.